OpenStudy (kainui):

Integrate: \(\int_0^\infty \frac{(\sin x)^{k}}{x^{k-n}\,dx\) for positive, odd \(k\) and \(0

mathslover (mathslover):

\(\Large{ \int_0^\infty \frac{( \sin x)^{k}} {x^{k-n}}\,dx } \)

OpenStudy (kainui):

Yeah, exactly haha. =P

OpenStudy (anonymous):

replace sinx with his taylor series

OpenStudy (anonymous):

its*